OpenStudy (anonymous):

Which improper fraction is equivalent to 3 2--- 9 14 -- 9 29 -- 9 21 -- 9 15 -- 9

OpenStudy (ash2326):

\[\large 2 \frac{3}{9}=2+ \frac{3}{9}=\frac{2 \times 9}{1 \times 9}+\frac{3}{9}=\frac{18}{9}+\frac{3}{9}=\frac{18+3}{9}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks

OpenStudy (ash2326):

@Arieboo11 did you understand?

OpenStudy (anonymous):

yes but witch one is it

OpenStudy (ash2326):

I won't tell you the last part is simple addition, do it and you'll get your answer

OpenStudy (anonymous):

thats the point of me asking i dont no math

OpenStudy (anonymous):

that good

OpenStudy (ash2326):

Just answer me what's 18+3? If you don't know this, I'll explain you and will help you understand:)

OpenStudy (anonymous):

21

OpenStudy (anonymous):

oh ok 21 over 9

OpenStudy (ash2326):

Good , You got it:D

OpenStudy (anonymous):

thank you so much

OpenStudy (ash2326):

@Arieboo11 You're welcome, if you ever need help here, do let me know. I'd be glad to help you:D

OpenStudy (anonymous):

k